Determine whether it can serve as the probability distribution of a random variable X. Explain your answer.

P(X)= 3+x / 3-x for X= 1, 2, 3, 4


1
Expert's answer
2022-01-18T11:48:07-0500

We are given that,

"P(X)= (3+x) / (3-x) \space for \space x= 1, 2, 3, 4"

Now,

"P(1)={(3+1)\over(3-1)}={4\over2}=2\\
P(2)={(3+2)\over(3-2)}={5\over1}=5\\
P(3)={(3+3)\over(3-3)}={6\over0}=\infin\\
P(4)={(3+4)\over(3-4)}={7\over-1}=-7\\"

For "P(x)" to be a probability distribution then it must satisfy the condition that "\sum P(X)=1"

So,

"\sum P(X)=2+5+\infin+-7=\infin"

Clearly, "P(X)" cannot serve as a probability distribution because, "\displaystyle\sum_{\forall x} P(X)\not=1"